ID |
Date |
Author |
Type |
Category |
Subject |
49
|
Fri May 2 21:39:13 2008 |
Dmass | Laser | PSL | Laser power on table lowered |
I have changed the polarizer before the polarizing beam splitter to minimize output power.
I did this to measure the PMC pole at as low a power as possible.
This will make power trends (which are picked off after the PBS) look VERY low for a bit. |
48
|
Fri May 2 21:09:26 2008 |
Dmass | Laser | General | Strange Noise |
We may want to set up a microphone and see if the noise is correlated to the sharp power fluctuations in the laser.
Quote: | Something on the table or under it is periodically making a squeeking noise. Stefan claims this is not new though neither of us can localize it (it's < 1sec duration). Unsure as to what is the offending noisemaker.
[Update - the noise happens when the laser is off, comes from side of table (possibly beneath) opposite laser] |
|
47
|
Fri Apr 25 13:33:35 2008 |
Dmass | Laser | PSL | power drift |
Here is a trend of the power drift taken over 10 days. The units are in counts produced by a photodiode stationed at a lower power pickoff of the beam.
The first plot is a full 10 days - you can see the times during which the laser was off where the counts go to zero.
The second plot is the full data over 10 hours, which seems to imply that the sharp increases in power are real.
Negative values correspond to an increase in power. |
Attachment 1: trend2.png
|
|
Attachment 2: trend3.png
|
|
46
|
Tue Apr 22 20:07:50 2008 |
Dmass | Laser | PSL | Profile of 35W PSL |
The total power seemed to be down about 10%, and that lower number was divided between two beams at the output. They opened up the amplifier and realigned things within it, which seems to have fixed the multibeam/lower power we were encountering before.
Quote: |
Quote: | I scanned the beam on the 35W laser again (for the first time since it was fixed by our German friends during the LSC meeting).
|
What did they fix? And how? |
|
45
|
Tue Apr 22 16:25:24 2008 |
rob | Laser | PSL | Profile of 35W PSL |
Quote: | I scanned the beam on the 35W laser again (for the first time since it was fixed by our German friends during the LSC meeting).
|
What did they fix? And how? |
44
|
Fri Apr 11 17:15:49 2008 |
Dmass | Laser | EOM | Transfer Function |
We traded network analyzers and now can go below 100kHz in our measurement. The following transfer function was taken with an HP 4395A.
The previous measurements were done with the PMC not locked to the laser, so the measurement was repeated using RF Reflection locking (sidebands @ 35MHz).
Another important thing we discovered was that the frame of the table is ungrounded, which led to significant noise leaking into the error signal when we let a BNC cable touch it. Schematic of experimental set up to follow.
Not all of the cables/electronics were calibrated out in the measurement below. |
Attachment 1: EOM_PMC_tf.png
|
|
43
|
Fri Apr 11 16:37:03 2008 |
Stefan Ballmer | Computing | General | Set up CDS wireless router / controlling from nokia handheld works |
I set up a CDS wireless router:
IP: 131.215.113.2
SSID: cdsrana
Unname & PWD written on router
It has WEP and MAC-filtering enabled.
With that I was able to use the little Nokia handheld to control the digital system.
Example:
- start x-terminal
- 'cd'
- 'ws1'
(this just starts ssh to ws1)
- startSlider
(this bring up the offset slider for the PZT control loop), or
- startmedm
(this starts the OMS main screen)
We have to make medm screens that work for the little Nokia screen.... |
42
|
Mon Apr 7 16:43:15 2008 |
Dmass | Laser | PSL | Profile of 35W PSL |
I scanned the beam on the 35W laser again (for the first time since it was fixed by our German friends during the LSC meeting).
A picture of the setup is below.
Images of the profile, 2d and 3d are below as .bmp, and the generating data for these is in Apr7_08scan.asc (zipped text)
This profile is significantly cleaner than the one measured in entry 31 of the elog. |
Attachment 1: scancap.png
|
|
Attachment 2: 3dbeamprof.bmp
|
Attachment 3: flatprof.bmp
|
Attachment 4: Apr7_08scan.asc.zip
|
41
|
Thu Apr 3 17:56:35 2008 |
Dmass | Laser | EOM | Transfer Function |
Transfer function part 2:
I repeated the measurement with impedence matching where I could (not inside the laser box), and did some CRUDE fitting with matlab.
I also herby acknowledge that png > jpg.
The tau is the time delay which makes the green curve on the phase plot coincide with the data at high frequency.
No locking was used. The 00 mode was reached for short periods of time through input to the laser crystal temperature during which time the measurement was taken.
-Calibration-
The input to the EOM was removed and connected to the PD output and calibrated out with the network analyzer.
The analyzer used to do this measurement has now been traded for a HP 4395A (which goes down to 10Hz). |
Attachment 1: EOMPMCTF.png
|
|
40
|
Tue Apr 1 20:50:37 2008 |
Dmass | Laser | EOM | Transfer Function |
(With Stefan's help) I took the transfer function between the EOM of the PSL, and the Photodiode on the PMC's reflected light. We looked at the lower frequency as well with another analyzer (up to 50 MHz).
The very low frequency features seem to be junk, and we can see what looks like ~f in the low frequency transfer function.
In the higher frequency range, there seem to be resonances corresponding to a FSR of about 22 MHz (long.) starting around 11 MHz. The low end seems to behave as expected, and go as frequency. (Note that the units on the non-log bottom plot are in x10^7 Hz) |
Attachment 1: EOM_PD_xsfer.jpg
|
|
Attachment 2: HFlogxsfer.jpg
|
|
Attachment 3: HFxsfer.jpg
|
|
39
|
Thu Mar 27 13:25:38 2008 |
Stefan Ballmer | Laser | ISS | Modifications to ISS (D080149-00-C) |
|
Attachment 1: D080149-00-C.pdf
|
|
38
|
Fri Mar 7 15:33:14 2008 |
Dmass | Laser | EOM | EOM shipped |
I packaged up and shipped the EOM back to UF. The translation stage, also theirs, was sent with it. Everything is marked c/o Guido Mueller. |
37
|
Tue Mar 4 16:00:28 2008 |
Dmass | Laser | PSL | ISS |
I moved the (non logging) ISS PD after the PMC, and changed the mirror/camera setup some.
Setup post PMC should now have ~5% of the power going to the camera, with the rest going to a BS for the ISS photodiodes. |
36
|
Mon Mar 3 01:21:19 2008 |
rana | Laser | MOPA | power drift |
This plot shows the power drift over the last few days.
The SIDE_OUT16 channel is the one. The y scale is zoomed in here but
zero corresponds to zero power. The faster spikes have a ~45 minute time scale.
No idea what this is yet. |
Attachment 1: test.png
|
|
Attachment 2: test2.png
|
|
35
|
Wed Feb 27 17:56:45 2008 |
Stefan Ballmer | Electronics | General | |
The 2nd HP 6209B 320V power supply initially didn't work because one of its backplane connectors had fallen off.
It has been fixed, and is now on the shelf, ready to be used. |
34
|
Wed Feb 27 17:14:47 2008 |
Stefan Ballmer | Electronics | General | PZT driver powered up |
- Powered up the PZT driver board using the 300V power supply (the one that's working).
- I haven't measured the noise or TF of the PZT driver box, but I
connected it to Ch3 (T3_ACT) and loaded a 10^2:1^2 anti-dewhitening filter.
(That's what the schematics asks for.)
- Removed the SR560 from the control loop & connected the demod error signal
to ch3 (TOP3).
- I put the control filter into the input filter module, because I needed the DOF4
control filter bank for a relief servo, and we need an offset adjust after the
relief servo for locking. |
33
|
Tue Feb 26 21:05:47 2008 |
Stefan Ballmer | Laser | General | |
Additionally to what Rana mentioned, there are two more signals hooked up:
SUS_LEFT (Ch4): PMC PZ control signal (from the SR560 loop)
SUS_RIGHT (Ch5): PMC transmitted power
We also hooked up the SUS_L_ACT (Ch4) DAC to the Laser temperature control.
Using this we closed the slow loop for the PMC |
32
|
Tue Feb 26 19:52:08 2008 |
Rana | Electronics | CDS | channels |
Stefan, Rana
We changed the cable hookup -- to check out the AI box + ADC.
Ch1 = 3 Vpp offset from a SRS function generator C2:OMS-SUS_TOP1_OUT_2048
Ch2 = 50 Ohms C2:OMS-SUS_TOP2_OUT_2048
Ch6 = ISS SENSE PD DC output C2:OMS-SUS_SIDE_OUT_2048
The laser is now on and set to record the power fluctuations over a long term. The value of the SIDE input gain
is set such that SUS_SIDE_OUT16 is in units of mW on the Ophir high power head.
--------------
We also noticed that the signals coming out of the AI box are very aliased. The thing is running at 2048 Hz but we think
that the AA and AI filters have 10 kHz cutoffs (D070081-00).
Next:
-----
- Redo the FE diagram to get rid of OMC SUS and make a PSL diagram.
- Get the trending working right.
- Up the FE sample rate to 32768 Hz
|
31
|
Sat Feb 16 15:33:25 2008 |
Dmass | Laser | PSL | Radial Beam Profile |
The following were taken with the beam scanner (having gone through a number of optical elements).
After leaving the main laser housing, the beam hits:
polarizer -> PBS -> low reflectivity pickoff mirror -> 1m lens -> 0.5m lens -> 45-P HR mirror -> 45 P HR mirror -> beamscan.
In taking my initial waist measurements, I had the beamscan after the LR pickoff, and the beam looked similar. |
Attachment 1: 2_16_0835w.bmp
|
30
|
Sat Feb 16 13:46:19 2008 |
Dmass | Laser | PSL | Laser Aperture closed. |
Laser was powered up with the aperture closed. It was fairly warm to the touch. I opened the aperture. I don't know if it is a "bad thing" to be dumping all 35W on the stop for large times. Sat @ 1:40. |
29
|
Tue Feb 12 17:00:35 2008 |
Dmass | Laser | PSL | Mode Matching/Lock |
I got a solution for mode matching the 35W to the PMC with 2 100mm (focal length) lenses. I abandoned this since they were deemed too fast by Stefan, and were seperated by ~1 focal length. We ended up using a 1m and .5m lens (all lenses from the newport kit in 058). My code choked on this solution and needs some bugs ironed out (will attach afterwards), so we used Stefan's.
We then managed to get a 00 lock, at a somewhat meager power throughput. Pictures of the layout below: |
Attachment 1: PMC_MM.JPG
|
|
Attachment 2: lenses_mmJPG.JPG
|
|
28
|
Tue Feb 12 16:54:13 2008 |
Dmass | Laser | PSL | 35W profile/waist |
The beam output by the 35W in 058 was significantly elliptical (up to a 10% difference in the waist measurements). I used the profile given by the "more gaussian" axis, and found a waist 9.7 cm inside the front of the enclosure, 250 microns in radius. We ended up using this axis to mode match to, but this ellipticity will probably need to be addressed as we try to maximize the power output through the PMC.
[plots to be added] |
27
|
Sun Feb 10 13:27:57 2008 |
Stefan Ballmer | Computing | General | Pictures of Bork-Space setup |
|
Attachment 1: CIMG3586.JPG
|
|
Attachment 2: CIMG3587.JPG
|
|
Attachment 3: CIMG3588.JPG
|
|
Attachment 4: CIMG3589.JPG
|
|
26
|
Sun Feb 10 13:18:07 2008 |
Stefan Ballmer | Laser | General | Trend of Laser Power over 20h |
Attached is a snapshot of the laser power over the last 24h.
The input gain is such that 1 count corresponds to about 1 mWatt
of power after the polarizing beam splitter, i.e. the the lambda/2 plate
was set to about 4 Watts.
I don't know where the short surges of power come from. |
Attachment 1: LaserPower20h.png
|
|
25
|
Tue Feb 5 18:59:07 2008 |
Stefan Ballmer | Laser | General | Bork-space running |
- Used the new Bork-space system to lock the PMC with a UGF of 100Hz.
- Hooked up a temporary power monitor (using the DC of one of the ISS PD's)
to ADC Channel 6. For the moment this can be used to trend the power.
- Accidentally blew the output opamp in the ISS monitoring chain - i.e. the ISS out-of-loop
monitor is currently unusable. Needs to be fixed. |
24
|
Wed Jan 30 12:57:53 2008 |
Dmass | Laser | PSL | Characterizing Beam |
1/28/08 - Took a series of beam waist measurements on the YAG.
-The polarizations (as measured by the photon beamscanner) had very different profiles. One polarization has a dual peak (probably a reflection from some optic). This shows up after the beam has passed through the polarizer, the PBS, and one mirror.
-Initial beam characterization yielded a waist inside the Laser Box by ~3", at a width of 200 microns. Will be rechecking this. (Impossible if there is a lens at the opening, which I'm not sure is the case). |
23
|
Tue Jan 29 12:50:00 2008 |
Dmass | Things to Buy | General | Things to buy wiki |
Temporarily located on the 40m wiki until AdhikariWiki is released.
http://lhocds.ligo-wa.caltech.edu:8000/40m/Adhikari_To_Buy |
22
|
Tue Jan 29 12:47:44 2008 |
Dmass | Laser | General | Strange Noise |
Something on the table or under it is periodically making a squeeking noise. Stefan claims this is not new though neither of us can localize it (it's < 1sec duration). Unsure as to what is the offending noisemaker.
[Update - the noise happens when the laser is off, comes from side of table (possibly beneath) opposite laser] |
21
|
Mon Jan 28 18:02:04 2008 |
Stefan Ballmer | Laser | General | 1.2MHz Laser intensity oscillation |
While working on the PDH loop for the PMC I noticed
a 1.2MHz oscillation on the laser intensity:
- It seems to be a bistable state: it was generally
started by some transient in the piezo input of the laser.
- The oscillation persists even when all control inputs to the laser
(EOM, PZT, AOM) are disconnected.
- It's extremely strong: about 20% fractional modulation
- Power-cycling the laser usually fixes the problem. |
20
|
Mon Jan 28 16:51:06 2008 |
Stefan Ballmer | Laser | PSL | PDH locked pre-mode cleaner |
Since we suspect that the ISS might be jitter-noise limited I started looking at the Pre-Mode Cleaner (PMC).
- Did a rough mode-matching with one lense into the PMC
- Aligned the PMC. Note that the camera is now in transmission of the PMC.
- Used 3MHz drive to the laser's EOM and a photo diode in reflection of the PMC for
PDH locking. Successfully closed the loop using a mixer and SR560, driving
the laser's piezo - however only on a 2-0 mode.
- The 30V power supply I currenctly use to pre-set the PMC PZT does not have
enough range to scan a full FSR, and I can't get the 00 mode into range.
David - feel free to dismantle the path into the PMC and properly measure the beam parameters. |
19
|
Thu Jan 24 19:43:53 2008 |
Stefan | Electronics | ISS | ISS servo work |
- Modified ISS: Bypassed differential driver & moved 2kHz pole to the first summing opamp.
- Installed an out-of-loop diode.
- Out-of-loop noise at 90Hz: ~8e-8 RIN/rtHz
Plot 1 shows the OLG.
The loop gain was set to 150kHz - just because I
didn't have the gain adjust hooked up.
Plot 2 shows the out-of-loop intesnity noise.
It is acoustic noise limited (except for the power lines). |
Attachment 1: ISSOLG.pdf
|
|
Attachment 2: ISSnoise.pdf
|
|
18
|
Wed Jan 2 15:10:24 2008 |
tobin | Lab Infrastructure | HVAC | Lab Temperature |
The lab temperature appears to have been stable over the holiday, always between 68.5 and 70.0 degrees F. (There is no data from Dec 28-29 due to a computer reboot.) |
17
|
Wed Dec 26 21:30:05 2007 |
rana | Laser | PSL | laser power |
I have rewired the laser power so now the Diode Box goes directly to one of the
overhead umbilicals (circuit #29) which is energized by the wall switch which
is marked as "Master Laser". Cutting off that switch is NOT the desired way to
power down the laser; use the Beckoff touchpad instead. The wall switch is the
emergency cutoff.
The HEPA fans on the laser table are both routed to power strips on the table now.
With the laser on, the beam looks well clipped. As far as I can tell this comes
from inside the box so the next thing is to do a little more re-alignment inside
the box.
Also need some more lenses for cameras and more IR cards.
And we should slice the tires of anyone who borrows equipment without leaving a note! |
16
|
Fri Dec 21 21:27:12 2007 |
tobin | Lab Infrastructure | HVAC | Lab Temperature |
Extech: 67.9°F, 26%RH. Met One: 68°F, 20%RH. Clear & cold outside.
Particle counter log looks stable. |
15
|
Tue Dec 18 13:49:50 2007 |
tobin | Lab Infrastructure | HVAC | Lab Temperature |
Extech: 68.6°F, 57%RH. Met One: 68°F, 53%RH. Drizzling outside.
Particle counter log shows stable temperature 67.5-69.5°F over the last four days. |
Attachment 1: temperature-20071218.pdf
|
|
14
|
Thu Dec 13 18:47:02 2007 |
tobin | Lab Infrastructure | HVAC | Lab Temperature |
Extech: 66.8°F, 25%RH. Met One: 68°F, 20%RH.
Here is the datalogger's contents:A 121307 161942 0100 0.3 000000 0.5 000000 TMP 000680 R/H 000130 LOC 000000 C/S 000E54
A 121307 164641 0100 0.3 000000 0.5 000000 TMP 000680 R/H 000130 LOC 000000 C/S 000E53
A 121307 171340 0100 0.3 000000 0.5 000000 TMP 000685 R/H 000130 LOC 000000 C/S 000E52
A 121307 174039 0100 0.3 000000 0.5 000000 TMP 000680 R/H 000200 LOC 000000 C/S 000E53
A 121307 180738 0100 0.3 000000 0.5 000000 TMP 000680 R/H 000210 LOC 000000 C/S 000E57
A 121307 183437 0100 0.3 000000 0.5 000000 TMP 000680 R/H 000210 LOC 000000 C/S 000E56
A 121307 190136 0100 0.3 000000 0.5 000000 TMP 000680 R/H 000195 LOC 000000 C/S 000E5C
A 121307 192835 0100 0.3 000000 0.5 000000 TMP 000685 R/H 000195 LOC 000000 C/S 000E69 I'm not sure why it only starts at 12:13pm today. Also, looks like its clock is 50 minutes fast. |
13
|
Wed Dec 12 16:19:00 2007 |
tobin | Lab Infrastructure | HVAC | Lab Temperature |
I connected the particle counter to the Dell computer for PEM data logging (temperature, humidity, and particle count). The data shows that while the temperature is comfortable right now, it was quite warm in here yesterday (78 degrees). A plot is attached.
For this I made a serial cable using two DB9-to-RJ45 converters and a length of Cat5 ethernet cable. The computer end of the cable has a female connector; the particle counter end of the cable has a male connector. Only three conductors are connected: transmit (pin 2), receive (pin 3), and ground (pin 7 on the particle connector, pin 5 on the computer). I used the USB-to-serial dongle, which shows up in Linux as /dev/ttyUSB0.
Someone later can write a script to automatically poll the particle counter for new data. For now, I just emptied its internal buffer, which stores the last 200 measurements. The secret to communicating with the particle counter, which I learned by examining the c1psl statecode at the 40m, is to first transmit the letter "U". After this, the particle counter will listen to you and do what you say. In particular, sending a capital letter "A" will cause the counter to emit the next measurement from its buffer, until there are none left, at which time it will just respond with "#". To read the data I just ran a "cat /dev/ttyUSB0" command in one terminal, while in another terminal I ran a loop "while 1; echo -n A > /dev/ttyUSB0; sleep 1; end". I set the serial port parameters (9600BPS, 8N1) using Minicom.
Also, there's a Matlab script to parse and plot the output, in ~controls/tobin. |
Attachment 1: temperature.pdf
|
|
12
|
Wed Dec 12 16:10:19 2007 |
tobin | Lab Infrastructure | HVAC | Lab Temperature |
The lab temperature at 2pm was a comfortable 69 degrees. |
11
|
Wed Nov 28 16:20:04 2007 |
Andrey | Lab Infrastructure | General | Enclosure cleaning |
Previous Tuesday, Nov. 20, I tried to wipe the curtain plastic enclosure around the optical table with the laser using methanol. It turned out, though, that methanol damages the surface of the enclosure, dissolving it and leaving lines/ traces. I stopped my activities with methanol after I realized that negative effect on three corner strips of enclosure.
Lesson for everyone: NEVER USE METHANOL FOR THOSE CURTAINS!
I tried 25% solution of isopropil alcohol in water on Wednesday, Nov. 21, and such mixture did not have any bad effect on the surface of the enclosure. I wiped the curtains, so they should be clean of dust now.
Andrey. |
10
|
Tue Nov 13 19:06:18 2007 |
rana | Laser | PSL | ISS servo work |
400 kHz !? Wow.
We should lower the resistors in the PD head so that we can put ~20-30 mA per diode
and use the ISS style diodes for both the in-the-loop as well as out-of-loop diodes.
Then we have to dump the beams reflecting off the diodes onto razor dumps and also
put in a lens so that the beam has a 0.3 mm dia. on the diodes.
Then we make an acrylic box with an entrance hole for the beam and put it over both
diodes. |
9
|
Mon Nov 12 21:33:41 2007 |
Stefan | Laser | PSL | ISS servo work |
I fixed a few things on the ISS board:
- Fixed the input THS4131 on the ISS board
- Fixed all monitoring point cables for the inner loop part.
With that I was able to lock the ISS with up to 400kHz UGF (limited by variable gain at max).
The sensing noise is about 60nV/rtHz at 100Hz without light, but there is some additional noise pick-up when the VCO driver is connected.
The In-loop noise, with 6V DC (~8mWatt on the diode) and a UGF of 400kHz is
100Hz: 170nV/rtHz (2.8e-8/rtHz)
200Hz: 60nV/rtHz (1.0e-8/rtHz)
500Hz: 50nV/rtHz (8.3e-9/rtHz)
The out-of loop noise is still a lot worse ( a few e-7/rtHz at 100Hz), but that's not surprising:
The table is quite noisy and not much care went into choosing good mounts. |
8
|
Thu Nov 8 18:54:41 2007 |
Stefan | Laser | PSL | 35W laser ISS loop closed with LIGO ISS electronics |
I closed the ISS loop using a iLIGO type ISS board, with an additional 2kHz pole at the last stage.
I haven't carefully measured noise or OLG yet, but it is not famous yet:
Looking at the in-loop diode I get a RIN of ~3e-7/rtHz at 100Hz, and ~1e-8 at 10kHz.
The gain is limited by two sharp resonances that rise at ~40kHz when the gain is increased.
TBD:
- Fix the ISS test point wiring - these cables have 800Ohm resistance!
(No not impedance... shield to core measures 800Ohms at DC...)
This brought back painful mamories from ~3years ago when we installed
the H1 ISS... apparently the known problem wasn't fixed on all versions.
- Measure OLG, in particular check whether weird AOM transfer function still
exists - it could be killing our gain.
- Measure sensing noise in-situ
- Install sensitive OL PD |
7
|
Thu Nov 8 01:58:37 2007 |
dmass | Laser | General | Tip Tilt Frequency Response |
Quote: | Here are screens of the Tip Tilt mirrors transfer function and associated xml files. Prettier unified plots will follow. The first one has the wrong header, the 071024 file is displayed with the 071014 header. |
Here are the plots with more useful units, and a doodle of the 'experimental' set-up at the end
Scroll down for a close up in the HF range.
The embedded .jpg is also available at the bottom of this entry as PitYaw.pdf |
Attachment 1: allfour.jpg
|
|
Attachment 2: TTstruc.JPG
|
|
Attachment 3: Pit_Yaw.pdf
|
|
6
|
Sun Nov 4 23:44:36 2007 |
rana | Laser | PSL | More Laser setup info |
Brought over a PMC mount from the 40m which Nergis is letting us borrow. Steve has made up a plate to
adapt that up to a 4" height from 3".
Still to do to get the optics of it set up:
- RF Amp to boost the 35.5 MHz Wenzel
- Attenuators to get the splitter output to not destroy the LO
- Bandpass filters for 35.5 MHz
- RFPD to take the PMC reflection and do locking
- 2 steering mirrors
- Mode matching lenses
|
5
|
Fri Oct 26 15:38:43 2007 |
Tobin Fricke | Laser | PSL | Polarizer |
On Tuesday we installed a λ/2 plate and a polarized beamsplitter after the laser aperture; attached to this entry is a measurement of transmitted power versus polarizer angle. |
Attachment 1: polarizer.pdf
|
|
Attachment 2: polarizer.m
|
% Polarizer calibration / Rana's lab
% Tobin Fricke 2007-10-26
% Experimental setup:
% [Dump]
% |
% +-------+ |
% | Laser |-------|lambda/2|----|PBS|----[Power Meter]
% +-------+
%
... 53 more lines ...
|
4
|
Wed Oct 24 23:50:27 2007 |
dmass | Laser | General | Tip Tilt Frequency Response |
Here are screens of the Tip Tilt mirrors transfer function and associated xml files. Prettier unified plots will follow. The first one has the wrong header, the 071024 file is displayed with the 071014 header. |
Attachment 1: TTL_low_freq.png
|
|
Attachment 2: TTL_mid_freq.png
|
|
Attachment 3: TTL_high_freq.png
|
|
Attachment 4: AngleTF_UL-071019.xml.zip
|
Attachment 5: AngleTF_UL-071020.xml.zip
|
Attachment 6: AngleTF_UL-071024.xml.zip
|
3
|
Mon Oct 22 20:04:28 2007 |
rana | Misc | General | Magnetic Levitation |
Links
Wikipedia article
Virgo Magnetic Suspension Idea from Monica
A paper by Giazotto on electrostatic mirrors is attached here.
And a document from Ron Drever on his magnetic levitation scheme:
here
Fender produces electric guitar pickups using Samarium Cobalt Noiseless (SCN) pickups. |
Attachment 1: Giazotto-ESmirror.pdf
|
|
2
|
Sun Oct 21 23:11:20 2007 |
rana | Laser | MOPA | Laser Turn on report |
When the laser was first turned on (slowly, as per the manual), the output power was only ~14 W instead of 35.
Some tweaking of the mirrors M4 & M5 (~1 full turn each in both pitch and yaw) restored the power. The peak
power was 37.0 W as measured on the Big Head. 
The beam is also not centered on the monitor PD and so its more likely a small shift of some earlier
mount than such large shifts of those last two mirrors.
With the cold temperature, the power has drooped to 35.5 W.  |
1
|
Sun Oct 21 23:06:49 2007 |
rana | Lab Infrastructure | HVAC | Its too cold again |
Friday evening it was a pleasant 70 deg F in the lab.
Tonight (Sunday evening) its back down to 61 deg F.
This happened ~a week ago; emailed Jeanne and Jo Ann and it got fixed. Will
try again and post results here. |
2705
|
Wed Dec 31 15:59:59 1969 |
Stephen | DailyProgress | Cryo vacuum chamber | Radiative Cooling of Si Mass, with worse inner shield inner surface emissivity - retry run was successful |
This post will host plots and trends from this radiative cooling run (QIL/2704).
Preliminarily, it looks like the reconfiguration to remove a hardware mistake or two led to a healthier run. The comparison below clarifies the two runs:
- QIL/2702 - conductive link between inner shield and outer shield (twisted pair from an RTD lead accidentally clamped); possibly another conductive link between outer shield and baseplate (outer shield more wobbly than usual on spacers)
- this data set should only be used to study the impact of a known conductive link between inner and outer shields.
- this run demonstrates that there will be more effective, faster cooling if the outer shield is conductively cooled!
- QIL/2704 - resolved above mistakes!
- this data set may be used to gain understanding of the impact of emissivity changes to the inner surface of the inner shield.
- may be compared to QIL/2695, a run that is equivalent except with a higher emissivity inner surface of the inner shield
Run ended with cryocooler shutdown at 12:27 pm (actual duration just under 92 hours). System will warm up with pumps on for the rest of the break, unless I am inspired to come in and run one of the next intended runs discussed in QIL/2704. I did not run any heat input test for this data set, as I am not planning to come in frequently enough to monitor the heating safely.
Data:
Attachment 1 compares QIL/2704 (solid) to QIL/2702 (dashed). As expected, the outer shield temperature from the latter run stays warm since the conductive short was resolved. Due to the reduction of the inner shield's thermal load, the inner shield is able to cool faster and plateau at a colder temperature. As Stephen pointed out, however, the test mass is not cooled as efficiently compared to when the outer shield was conductively cooled.
Fitting Results:
Attachment 2 is a current model diagram of the various components being considered, and their thermal couplings. Attachment 3 plots the fitted model (dashed) over the temperature data (solid). The fit parameters were the following emissivities: aluminum foil, rough aluminum, and aquadag. Notes from the fit:
1. With the conductive shorting of the outer shield resolved, the model (which considers only radiative cooling of the OS) is well fit to the OS temperature data.
2. The inner shield model is missing some key term(s) affecting its time constant and steady state temperature.
3. The above error propagates to the test mass model (I believe).
Given these caveats, the fit results are as follows: aquadag e = 0.92, Al foil e = 0.04, rough Al e = 0.19. These all initially seem reasonable, and I'm happy to see that the aquadag emissivity is higher than previously estimated.
Next steps:
1. Separate the cold plate from the inner shield, and model their conductive and radiative link. Also model the radiative link between the cold plate and the test mass.
2. Cover the test mass in foil (to best of our ability) to refine the radiative link between the test mass and inner shield. Doing so will mean both elements have the same emissivity, so there is only one unknown parameter. |
Attachment 1: cooldown_12-21_vs_12-10.pdf
|
|
Attachment 2: Megastat_Heat_Load_Sketch.png
|
|
Attachment 3: 12_21_cooldown_fit.png
|
|